Given Utility function U=20Q-2Q^2what quantity of good gives highest utility for the consumer? What will happened if the consumer obtains more quantity than the one you computed? Explain briefly


1
Expert's answer
2021-05-04T07:24:55-0400

Given utility function-

"U=20Q-2Q^2"


"\Rightarrow \dfrac{dU}{dQ}=20-4Q"


Foe Maximum quantity of goods-

Putting "\dfrac{dU}{dQ}=0\Rightarrow 20-4Q=0\Rightarrow Q=5"

The 5 quantity of goods gives the highest utility for the consumer.


When the consumer consume more quantity than calculate, the consumer equillibrium gets disturbed. The marginal utility decreases and the average utility diminishing and the consumer may suffer a loss.